OpenStudy (wkep1234):

Triangle ABC has side lengths 12, 16, and 20. Do the side lengths form a Pythagorean triple? Explain.

OpenStudy (lapinblanc):

Yes A pythagorean triple is three positive integars a, b, c such that a^2+b^2=c^2. Common example is (3, 4, 5) then so is (ka, kb, kc). In your example, you have 12, 16, 20 which would look like this: (12)^2+(16)^2=(20)^2 but if you break down your a, b, c values and see if they are multiples of 3, 4, 5 you find: (4*3)^2+(4*4)^2=(4*5)^2 Bc 4*3=12 or your a 4*4=16 or your b 4*5=20 or your c and k is 4 (ka, kb, kc) (k3, k4, k5) (4*3, 4*4, 4*5) (12, 16, 20)